# LOSC: LiDAR Open-voc Segmentation Consolidator

### v1

@@ Appendix A Results with super-classes
nuScenes SemanticKITTI
[…]
@@
Appendix B Classwise results
@@
These tables detail the results reported in Table 3 of the main paper. They also provide classwise results for super-classes (cf. Appendix A).
@@
For nuScenes, while the IoUof most classes keeps on increasing as we iterate more, only a few classes see their performance stagnate or slightly drop. Therefore, the overall performance (mIoU%) is clearly increasing even after 3 iterations.
@@
Appendix C Quantity and quality of labeled points for training
@@
Appendix D Text prompts
@@
Appendix E Augmentations
The following Table 15 presents detailed classwise results of zero-shot OpenSeeD [19] applied to the nuScenes dataset, before and after using our set of 10 distinct data augmentation methods. We observe that the results remain largely consistent with the original, which helps us eliminate obviously incorrect labels. We also present the results of several other data augmentations in Table 15. These augmentations perform significantly worse than the no-augmentation baseline and result in a substantial loss of points.
@@ Appendix E Augmentations
Appendix F Qualitative examples
Last, we present in Figure 3 qualitative results on the nuScenes and SemanticKITTI validation sets. A typical error with LOSC on both datasets is a confusion between different types of flat surfaces such as road/driveable surface, sidewalk and terrain. In SemanticKITTI, trunks are also systematically included in vegetation rather than considered as a separate class. Theses observations are consistent with the quantitative results provided in Tables 8 and 9.
Ground truth LOSC
nuScenes
Refer to caption Refer to caption
nuScenes
Refer to caption Refer to caption
SemanticKITTI
Refer to caption Refer to caption
SemanticKITTI
Refer to caption Refer to caption
Figure 3: Qualitative results of semantic segmentation from the validation sets of nuScenes and SemanticKITTI. The color code used to represent each class is provided in Fig. 4.
@@ Appendix F Qualitative examples
Figure 4: Color code used to represent each class on nuScenes (top) and SemanticKIITI (bottom) in Fig. 3.

### v2

@@ Appendix B Results with super-classes
Method $_{\text{sem}} mIoU$$_{\text{pan}}$ mIoU $_{\text{sem}}^{\text{\,(sc)}}$ mIoU $_{\text{pan}}^{\text{\,(sc)}}$ mIoU mIoU $^{\smash{\text{\,(sc)}}}$
@@
Appendix C Classwise results
@@
These tables detail the results reported in Table 3 of the main paper. They also provide classwise results for super-classes (cf. Appendix B).
@@
For nuScenes, while the IoU% of most classes keeps on increasing as we iterate more, only a few classes see their performance stagnate or slightly drop. Therefore, the overall performance (mIoU%) is clearly increasing even after 3 iterations.
@@
Appendix D Quantity and quality of labeled points for training
@@
Appendix E Text prompts
@@
Appendix F Augmentations
The following Table 15 presents detailed classwise results of zero-shot OpenSeeD [33] applied to the nuScenes dataset, before and after using our set of 10 distinct data augmentation methods from the Albumentations library [3]. We observe that the results remain largely consistent with the original, which helps us eliminate obviously incorrect labels. We also present the results of several other data augmentations in Table 15. These augmentations perform significantly worse than the no-augmentation baseline and result in a substantial loss of points.
@@ Appendix F Augmentations
Figure 4: Color code used to represent each class on nuScenes (top) and SemanticKIITI (bottom) in Figure 3.